Excel Online Structured Activity: NPV profiles A company is considering two mutually exclusive expansion plans. Plan A requirA company has a 13% WACC and is considering two mutually exclusive investments (that cannot be repeated) with the following cd. From your answers to parts a-c, which project would be selected? If the WACC was 18%, which project would be selected? e.
Answer #1

Answer 1

NPV is given by:

Σ Rt/(1 i)ttvaries from 1 to n NPV where Rt Cash flow netted (Inflow - Outflow) during the period t i Discount rate t numbe

The Discount rate at which NPV = 0 is known as IRR

Answer a)

X V fx IRR(C3:C23) SUM А В Discount Rate Year 9% 1 Plan B Plan A 2 -$39.00 -$13.00 3 $6.23 $2.91 4 1 $6.23 $2.91 5 2 $6.23 $2

X V fx =NPV(C1,C4:C23)+C3 SUM А В C E Discount Rate 9% Plan A Plan B 2 Year -$39.00 -$13.00 3 $6.23 $2.91 4 1 $6.23 $2.91 5 2

X V fx =NPV(C1,D4:D23)+D3 SUM В E 9% Plan B -$13.00 Discount Rate Year 1 Plan A 2 -$39.00 3 0 $6.23 $2.91 4 1 $6.23 $2.91 5 2

fx NPV(C1,D4:D23)+D3 D25 А В C E Discount Rate 9% Plan A Plan B 2 Year -$39.00 -$13.00 3 0 $6.23 $2.91 4 $6.23 $2.91 5 2 $6.2

-----------------------------

Answer b)

NPV profile

X V fx =NPV(F5,$C$6: $c$25)+$C$5 SUM D. A В E Н 1 2 3 Discount Rate NPV of plan A NPV of plan B 0%=NPV(F5, $C$6:$C$25 )+ $C$5

X V fxNPV(F5,$ D$6:$D$25)+$D$5 SUM В D E G 2 3 Year Discount Rate NPV of plan A NPV of plan B $85.60 NPV(F5, $D$6:$D$25)+ $D$

Discount Rate NPV of plan A NPV of plan B
0% $85.60 $45.20
1% $73.42 $39.51
2% $62.87 $34.58
3% $53.69 $30.29
4% $45.67 $26.55
5% $38.64 $23.27
6% $32.46 $20.38
7% $27.00 $17.83
8% $22.17 $15.57
9% $17.87 $13.56
10% $14.04 $11.77
11% $10.61 $10.17
12% $7.53 $8.74
13% $4.76 $7.44
14% $2.26 $6.27
15% -$0.00 $5.21
16% -$2.06 $4.25
17% -$3.94 $3.38
18% -$5.65 $2.58
19% -$7.22 $1.84
20% -$8.66 $1.17
21% -$9.99 $0.55
22% -$11.21 -$0.02
23% -$12.34 -$0.55
24% -$13.39 -$1.04
25% -$14.37 -$1.49
26% -$15.27 -$1.92
27% -$16.12 -$2.31
28% -$16.91 -$2.68
29% -$17.65 -$3.03
30% -$18.34 -$3.35
31% -$18.99 -$3.66
32% -$19.61 -$3.94
33% -$20.18 -$4.21
34% -$20.73 -$4.47
35% -$21.24 -$4.71
36% -$21.73 -$4.93
37% -$22.19 -$5.15
38% -$22.63 -$5.35
39% -$23.05 -$5.55
40% -$23.44 -$5.73
41% -$23.82 -$5.91
42% -$24.18 -$6.08
43% -$24.52 -$6.24
44% -$24.85 -$6.39
45% -$25.16 -$6.54
46% -$25.46 -$6.68
47% -$25.75 -$6.81
48% -$26.03 -$6.94
49% -$26.29 -$7.06
50% -$26.54 -$7.18
51% -$26.79 -$7.30
52% -$27.02 -$7.41
53% -$27.25 -$7.51
54% -$27.47 -$7.61
55% -$27.67 -$7.71
56% -$27.88 -$7.80
57% -$28.07 -$7.90
58% -$28.26 -$7.98
59% -$28.44 -$8.07
60% -$28.62 -$8.15
61% -$28.79 -$8.23
62% -$28.95 -$8.31
63% -$29.11 -$8.38
64% -$29.27 -$8.45
65% -$29.42 -$8.52
66% -$29.56 -$8.59
67% -$29.70 -$8.66
68% -$29.84 -$8.72
69% -$29.97 -$8.78
70% -$30.10 -$8.84
71% -$30.23 -$8.90
72% -$30.35 -$8.96
73% -$30.47 -$9.01
74% -$30.58 -$9.07
75% -$30.69 -$9.12
76% -$30.80 -$9.17
77% -$30.91 -$9.22
78% -$31.01 -$9.27
79% -$31.11 -$9.32
80% -$31.21 -$9.36
81% -$31.31 -$9.41
82% -$31.40 -$9.45
83% -$31.49 -$9.49
84% -$31.58 -$9.54
85% -$31.67 -$9.58
86% -$31.76 -$9.62
87% -$31.84 -$9.66
88% -$31.92 -$9.69
89% -$32.00 -$9.73
90% -$32.08 -$9.77
91% -$32.15 -$9.80
92% -$32.23 -$9.84
93% -$32.30 -$9.87
94% -$32.37 -$9.90
95% -$32.44 -$9.94
96% -$32.51 -$9.97
97% -$32.58 -$10.00
98% -$32.64 -$10.03
99% -$32.71 -$10.06
100% -$32.77 -$10.09

NPV Profile $140.00 $120.00 $100.00 $80.00 $60.00 $40.00 $20.00 Plan B $0.00 Plan A -$20.00 -$40.00 -$60.00 Discount Rate %00

Hence, cross over rate is near 11%

-----------------------------------------

Answer c)

For crossover rate

Step 1: Subtract cash flows of 2 project

Step 2: Find IRR of difference

X V fx =D3-C3 SUM A B C E F Discount Rate 9% 1 Plan A Plan B Plan A - Plan B 2 Year -$39.00 -$13.00 0 D3- C3 3 $6.23 $2.91 -$

A A = EE Cut Calibri Percenta 11 В IU Merge and Wrap Copy Format Painter Paste Center Тext X V fx =IRR(F3:F23) SUM A В C D F

fxIRR(F3:F23) F24 А В C E F G Discount Rate 9% Plan A Plan B Plan A Plan B 2 Year -$39.00 -$13.00 $26.00 3 0 $6.23 $2.91 -$3.

-----------------------------